[Brahms-l] Shift report 20040213 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Fri Feb 13 2004 - 15:57:29 EST
------------------------------------------------------------
Date: Fri Feb 13 15:57:28 2004, Shift: 08:00-16:00
Supervisor: EBJ, with: 
Log book#15 pages 55 - 55

Shift Summary:
------------------------------------------------------------
1.  The machine has been performing very well.  We had useful beam most of the shift.

2. I collected close to 400k trig 2 events at full field.
 The D2 magnet seems to be operating fine.  It has not tripped since I ramped it to full field around 9:30am.  Ramiro came in and looked at the hall probe for the D2 magnet at 1/4 field.  The current fluctuations in the current at 1/4 field do not show up in the hall probe.

3. No new detector problems.   
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Feb 13 15:57:43 2004

This archive was generated by hypermail 2.1.8 : Fri Feb 13 2004 - 15:58:08 EST